OpenStudy (anonymous):

How do you solve: 7 to the square root of 72? Please show step by step.

jimthompson5910 (jim_thompson5910):

Is the problem \[\large 7\sqrt{72}\]???

OpenStudy (anonymous):

Yes, please express answer in the lowest radical form with no decimal.

jimthompson5910 (jim_thompson5910):

\[\large 7\sqrt{72}\] \[\large 7\sqrt{36\times 2}\] \[\large 7\sqrt{36}\times \sqrt{2}\] \[\large 7\times 6\times \sqrt{2}\] \[\large 42\sqrt{2}\] So \[\large 7\sqrt{72}=42\sqrt{2}\]